Suppose that, on average, in every two pages of a book there is one typographical error, and that the number of typographical errors on a single page of the book is a Poisson r.v. with λ = 1/2. What is the probability of at least one error on a certain page of the book? Interpret your result.
Thank you

Answer by Boreal(15235)

Find the probability there are 0 errors on a page, and 1- that probability is the answer,
this is p(0)=e^(-0.5)*0.5^0/0!=e^(-0.5)=0.6065 probability
The answer is 1-0.6065 or 0.3935